# Natural Metric-Affine Inflation

### v2

@@ 2 The action and the dynamical degrees of freedom
$S_{\rm NI}=\int d^{4}x\sqrt{-g}\left[\alpha R-\frac{\partial_{\mu}\phi\,\partial^{\mu}\phi}{2}+\frac{\frac{\alpha}{4}\partial_{\mu}\alpha\partial^{\mu}\alpha-\alpha\partial_{\mu}\beta\partial^{\mu}\beta-2\beta\partial_{\mu}\alpha\partial^{\mu}\beta}{\frac{2}{3}\left(\beta^{2}+\frac{\alpha^{2}}{4}\right)}-V\right],$ (2.8)
@@
$\left.+\frac{\frac{\partial_{\mu}\alpha\partial^{\mu}\alpha}{8}-\frac{\partial_{\mu}\beta\partial^{\mu}\beta}{2}-\frac{\beta}{\alpha}\partial_{\mu}\alpha\partial^{\mu}\beta}{\frac{2}{3M_{P}^{2}}\left(\beta^{2}+\frac{\alpha^{2}}{4}\right)}-\frac{M_{P}^{4}V}{4\alpha^{2}}\right].$
@@
$k(\phi)$$\equiv$ $\frac{M_{P}^{2}}{2}\left[\frac{1}{\alpha}+\frac{3\alpha^{\prime 2}}{\alpha^{2}}\left(1-\frac{\alpha^{2}}{\alpha^{2}+4\beta^{2}}\right)+\frac{12\beta^{\prime 2}}{\alpha^{2}+4\beta^{2}}+\frac{24\beta\alpha^{\prime}\beta^{\prime}}{\alpha(\alpha^{2}+4\beta^{2})}\right]$ (2.12)
$=$ $\frac{M_{P}^{2}}{2}\left[\frac{1}{\alpha}+\frac{12(\alpha^{\prime}\beta+\alpha\beta^{\prime})^{2}}{\alpha^{2}(\alpha^{2}+4\beta^{2})}\right]>0,$ (2.13)
@@ 4.3 $\xi>0$ and $\tilde{\xi}>0$
Moreover, this case is interesting also because it allows us to satisfy the experimental constraints even when $f$ is sub-Planckian. The corresponding results are shown in Figs. 14 and 15. For numerical convenience the study considers only values of $\tilde{\xi}$ between approximately 10 and 20. The lines representing to the results corresponding to $\delta_{f}=10^{-2}$ and $\delta_{f}=10^{-1}$ are essentially undistinguishable apart for the $\alpha_{s}$ vs. $n_{s}$ plot. We can see that when $\xi=0$, with a large enough $\tilde{\xi}$, the $r$ vs. $n_{s}$ predictions at very low $\delta_{f}$ are again compatible with data. However, in this case $\alpha_{s}$ is very large and essentially always out from the allowed region for sub-Planckian $f$. This was somehow expected because in standard natural inflation the second slow-roll parameter $\eta$ is very large in absolute value at low $\delta_{f}$ (potentially even breaking the slow-roll approximation), implying a very small $n_{s}$. In order to achieve compatibility with data, $n_{s}$ is varying very fast causing its running $\alpha_{s}$ to be out of the allowed region. The issue can be solved by using $\xi>0$, which restores compatibility with data by moving the large $\alpha_{s}$ region towards smaller $n_{s}$ values.
@@ Acknowledgments
The work of A.R. was supported by the Estonian Research Council grants PRG1055, RVTT3, RVTT7 and the CoE program TK202 “Fundamental Universe”. The work of A.S. was also supported by the grant DyConn from the University of Rome Tor Vergata. This article is based upon work from COST Actions COSMIC WISPers CA21106 and CosmoVerse CA21136, supported by COST (European Cooperation in Science and Technology).

### v3

@@ 2 The action and the dynamical degrees of freedom
$S_{\rm NI}=\int d^{4}x\sqrt{-g}\left[\alpha R-\frac{\partial_{\mu}\phi\,\partial^{\mu}\phi}{2}+\frac{\frac{\alpha}{4}\partial_{\mu}\alpha\partial^{\mu}\alpha-\alpha\partial_{\mu}\beta\partial^{\mu}\beta+2\beta\partial_{\mu}\alpha\partial^{\mu}\beta}{\frac{2}{3}\left(\beta^{2}+\frac{\alpha^{2}}{4}\right)}-V\right],$ (2.8)
@@
$\left.+\frac{\frac{\partial_{\mu}\alpha\partial^{\mu}\alpha}{8}-\frac{\partial_{\mu}\beta\partial^{\mu}\beta}{2}+\frac{\beta}{\alpha}\partial_{\mu}\alpha\partial^{\mu}\beta}{\frac{2}{3M_{P}^{2}}\left(\beta^{2}+\frac{\alpha^{2}}{4}\right)}-\frac{M_{P}^{4}V}{4\alpha^{2}}\right].$
@@
$k(\phi)$$\equiv$ $\frac{M_{P}^{2}}{2}\left[\frac{1}{\alpha}+\frac{3\alpha^{\prime 2}}{\alpha^{2}}\left(1-\frac{\alpha^{2}}{\alpha^{2}+4\beta^{2}}\right)+\frac{12\beta^{\prime 2}}{\alpha^{2}+4\beta^{2}}-\frac{24\beta\alpha^{\prime}\beta^{\prime}}{\alpha(\alpha^{2}+4\beta^{2})}\right]$ (2.12)
$=$ $\frac{M_{P}^{2}}{2}\left[\frac{1}{\alpha}+\frac{12(\alpha^{\prime}\beta-\alpha\beta^{\prime})^{2}}{\alpha^{2}(\alpha^{2}+4\beta^{2})}\right]>0,$ (2.13)
@@ 4.3 $\xi>0$ and $\tilde{\xi}>0$
Moreover, this case is interesting also because it allows us to satisfy the experimental constraints even when $f$ is sub-Planckian. The corresponding results are shown in Figs. 14 and 15. For numerical convenience the study considers only values of $\tilde{\xi}$ between approximately 10 and 25. The lines representing to the results corresponding to $\delta_{f}=10^{-2}$ and $\delta_{f}=10^{-1}$ are essentially undistinguishable apart for the $\alpha_{s}$ vs. $n_{s}$ plot. We can see that when $\xi=0$, with a large enough $\tilde{\xi}$, the $r$ vs. $n_{s}$ predictions at very low $\delta_{f}$ are again compatible with data. However, in this case $\alpha_{s}$ is very large and essentially always out from the allowed region for sub-Planckian $f$. This was somehow expected because in standard natural inflation the second slow-roll parameter $\eta$ is very large in absolute value at low $\delta_{f}$ (potentially even breaking the slow-roll approximation), implying a very small $n_{s}$. In order to achieve compatibility with data, $n_{s}$ is varying very fast causing its running $\alpha_{s}$ to be out of the allowed region. The issue can be solved by using $\xi>0$, which restores compatibility with data by moving the large $\alpha_{s}$ region towards smaller $n_{s}$ values.
@@ Acknowledgments
A.R. thanks I.D. Gialamas for useful discussions. The work of A.R. was supported by the Estonian Research Council grants PRG1055, RVTT3, RVTT7 and the CoE program TK202 “Fundamental Universe”. The work of A.S. was also supported by the grant DyConn from the University of Rome Tor Vergata. This article is based upon work from COST Actions COSMIC WISPers CA21106 and CosmoVerse CA21136, supported by COST (European Cooperation in Science and Technology).